# Discovering the Signal Subgraph: An Iterative Screening Approach on Graphs

### v2

@@ Abstract
Supervised learning on graphs is a challenging task due to the high dimensionality and inherent structural dependencies in the data, where each edge depends on a pair of vertices. Existing conventional methods designed for Euclidean data do not account for this graph dependency structure. To address this issue, this paper proposes an iterative vertex screening method to identify the signal subgraph that is most informative for the given graph attributes. The method screens the rows and columns of the adjacency matrix concurrently and stops when the resulting distance correlation is maximized. We establish the theoretical foundation of our method by proving that it estimates the true signal subgraph with high probability. Additionally, we establish the convergence rate of classification error under the Erdos-Renyi random graph model and prove that the subsequent classification can be asymptotically optimal, outperforming the entire graph under high-dimensional conditions. Our method is evaluated on various simulated datasets and real-world human and murine graphs derived from functional and structural magnetic resonance images. The results demonstrate its excellent performance in estimating the ground-truth signal subgraph and achieving superior classification accuracy.
Keywords: iterative screening, distance correlation, graph classification
@@ 1 Introduction
The analysis of graph structure is critical in various big data fields, including neuroscience, internet mapping, and social networks (Otte and Rousseau, 2002; Newman et al., 2002; Bullmore and Bassett, 2011; Vogelstein et al., 2013; Shen et al., 2017; Lee et al., 2019). Due to the large size of graphs in practice, such as in social networks and raw neuroimages, it is often necessary to use smaller subgraphs from the observed graphs. Moreover, the selected subgraph should maintain or improve subsequent inference. For example, identifying a subset of brain regions from brain imaging to better predict the phenotype of interest in each subject.
The statistical problem of feature reduction and dimension selection has been extensively studied, with well-known methods such as Lasso (Tibshirani, 1996), adaptive Lasso (Zou and Hastie, 2006), Dantzig selector (Candes and Tao, 2007), sure independence screening (Fan and Lv, 2008; Li et al., 2012), among others. These methods have specific objectives, such as sparsity and the recovery of ground-truth. Among them, the screening method is known for its computational efficiency and model-free nature and are commonly used for high-dimensional data Zhu et al. (2011), making it a suitable candidate for graph data.
However, dimension reduction for graph data presents unique challenges because of its high-dimensionality and the unique structure of the $n\times n$ adjacency matrix. To that end, this paper proposes an novel iterative screening method on graph data, which utilizes distance-based correlation and independence screening in an iterative manner to estimate the signal subgraph. During each iteration, we define the feature for each vertex using the adjacency of the reduced graphs, compute a distance-based correlation between the feature and the label of interest $Y$, and discard vertices with small correlations. This process is repeated recursively on the reduced graphs from previous iterations, yielding a smaller set of vertices each time until an estimated signal subgraph is selected for output.
@@ 2.1 Setting and Notations
In the classical statistical pattern recognition setting, the pairs of observations $\{(A_{i},Y_{i})\}_{i=1}^{m}$ are independent and identically distributed pairs according to a distribution $F_{A,Y}$ (Devroye et al., 2013), that is
@@ 2.2 Distance Correlation
The distance correlation is a measure that can detect all types of dependencies between two random variables, given sufficient sample size (Székely et al., 2007). To compute the sample distance correlation, two pairwise distance matrices are transformed and multiplied using a Hadamard product. The sample distance correlation is asymptotically 00 if and only if the two underlying random variables are independent. For more detailed mathematical information about the distance correlation and its population definition, see the Appendix.
The distance correlation is a computationally efficient method (Shen et al., 2022), and has been shown to be equivalent to kernel correlation Shen and Vogelstein (2021). It has been used for various inference tasks (Wang et al., 2015; Fokianos and Pitsillou, 2018; Shen and Dong, 2024; Shen et al., 2023) not limited to screening. This paper also utilizes a local version of the distance correlation called the multiscale graph correlation (MGC), which improves testing power against nonlinear dependencies (Vogelstein et al., 2019; Shen et al., 2020).
@@ 4 Theoretical Properties
  • The graph adjacency matrix $A$ follows the Erdos-Renyi random graph model (Erdos and Renyi, 1959) and has a bounded Frobenius norm.
@@
The results suggest that using the estimated signal subgraph via iterative screening can be expected to perform better than using the whole graph, when the size of the signal subgraph is fixed and the number of observed graphs is comparable to the size of the whole graph. This setting is illustrated in the top panel of Figure 2 with $n=200$, $m=300$, and $|S|=20$, and is also observed in the experiments in Section 5 where the estimated subgraph yields better classification performance. All proofs and necessary mathematical background are provided in the appendix.
@@ 5.1 Signal Subgraph Estimation
We estimate the subgraph using various screening methods, including the conventional screening with Dcor and MGC, iterative screening with Dcor and MGC at $\delta=0.5$ and $\delta=0.05$ respectively, and screening with canonical correlation analysis (CCA) (Hotelling, 1936) and RV coefficient (RV) (Robert and Escoufier, 1976). Since the actual size of the signal subgraph was known to be 20, we output the estimated subgraph at the same size and calculated the true positive rate. The ROC curve is shown in Figure 1, while Table 1 reports the AUC and runtime for each approach. We observe that Dcor and MGC outperform CCA and RV, and terative screening improves the performance over conventional screening. Furthermore, iterative screening with $\delta=0.05$ yields better results than iterative screening with $\delta=0.5$ at the cost of a longer running time.
@@ 6.1 Site and Sex Prediction With Human Brain
Our objective is to predict the sex and site of each individual based on functional magnetic resonance image (fMRI) graphs (Ogawa et al., 1990). We utilized two datasets, SWU4 (Liu et al., 2017) and HNU1 (Chen et al., 2015), which include $467$ and $300$ subjects, respectively. Each individual’s fMRI scan is registered to the MNI152 template using the Desikan atlas, which has $70$ regions (Desikan et al., 2006). The graphs are created using the NeuroData’s MRI Graphs pipeline [^1] https://github.com/neurodata/ndmg, a popular tool for processing and representing brain images.
@@ 6.2 Sex Difference in Mouse Brain
Structural magnetic resonance imaging has provided insight into the genetic basis of mouse brain variability by examining the relationship between volume covariance and genotypes (Badea et al., 2009). With high-resolution diffusion tensor imaging and tractography, we can now investigate the underlying bases for structural connectivity patterns (Calabrese et al., 2015), in relationship with genotype and sex. Based on MRI and conventional Nissl histology, we scanned and registered $55$ mouse brains (pooled genotypes) into the space of a minimum deformation template, aligned to Waxholm space (Johnson et al., 2010). The atlas labels were propagated onto the template and, subsequently, onto each individual brain using ANTs (Avants et al., 2011). We employed DSI Studio (Yeh et al., 2013) to estimate tract-based structural connectivity for each brain, which was then represented as a graph with $332$ vertices, $166$ per hemisphere. Of the $55$ mice, $32$ are male, and $23$ are female.
Similarly, we conduct a leave-one-out evaluation using an iterative vertex screening to estimate the signal subgraph, followed by a $9$-nearest-neighbor classifier to predict the left-out sample based on the estimated signal subgraph. Figure 6 demonstrates the prediction error and distance correlation when using the iterative screening algorithm. Despite the small sample size and fluctuating prediction error, the screening method outputs a signal subgraph of size $10$, which results in a near-optimal classification error of $0.18$. The estimated signal vertices include a thalamic component and the periaqueductal gray, which play an important role in driving sexually dimorphic mouse brain development (Spring et al., 2007; Raznahan et al., 2013).
@@ 7 Conclusion
In summary, we developed an iterative vertex screening methodology to estimate the signal subgraph of interest, and apply the algorithm to analyze the MRI brain images. Utilizing distance correlation and multiscale graph correlation leads to a better measure on arbitrary dependency between the vertex and response variable. The experiments and theories offer strong evidence that the iterative approach estimates the signal subgraph effectively and accurately, which leads to better performance for subsequent inference task than conventional screening.
@@ Acknowledgement
8 More Preliminaries
8.1 Distance Correlation
Given sufficient sample size, the distance correlation (Székely et al., 2007) is able to detect all types of dependencies between two random variables. The population distance covariance $Dcov(X,Y)$ can be defined via either the characteristic functions or Euclidean distance as:
@@ 8.1 Distance Correlation
8.2 Graph Classification
We introduce the binary classification setting of predicting the label $Y\in\{0,1\}$ using graph $A$. This set-up serves as the basis for Section 4 and the simulations. The network model under consideration is the inhomogeneous Erdos-Renyi (ER) random graph model (Erdos and Renyi, 1959), which allows edges to have different probabilities and generates a family of distributions on undirected graphs. The ER model can also be viewed as a stochastic block model with each block containing only one vertex.
@@ 8.2 Graph Classification
Given this model, the optimal classification performance is achieved by the Bayes classifier $g^{*}(\cdot)$ (Devroye et al., 2013) defined as
@@
9 Proof
9.1 Proof of Theorem 1
@@
We have now met the two requirements to apply Theorem 1 in (Li et al., 2012). By utilizing the theorem and choosing $\kappa=0$, we can conclude that there exist two positive constants $c_{1},c_{2}$, and for any $20<\gamma<1/$ 2 we have
@@ 9.2 Proof of Theorem 2
Setting $2(e_{V}\epsilon_{2}+\epsilon_{1})=\epsilon$ and $2\epsilon_{1}^{2}=c_{4}\epsilon_{2}^{2}$, we have $\epsilon_{2}=\frac{\epsilon}{2e_{V}+\sqrt{2c_{4}}}$. Applying Theorem 2.3 in (Devroye et al., 2013) yields

### v3

@@ Abstract
Supervised learning on graphs is a challenging task due to the high dimensionality and inherent structural dependencies in the data, where each edge depends on a pair of vertices. Existing conventional methods are designed for standard Euclidean data and do not account for the structural information inherent in graphs. In this paper, we propose an iterative vertex screening method to achieve dimension reduction across multiple graph datasets with matched vertex sets and associated graph attributes. Our method aims to identify a signal subgraph to provide a more concise representation of the full graphs, potentially benefiting subsequent vertex classification tasks. The method screens the rows and columns of the adjacency matrix concurrently and stops when the resulting distance correlation is maximized. We establish the theoretical foundation of our method by proving that it estimates the true signal subgraph with high probability. Additionally, we establish the convergence rate of classification error under the Erdos-Renyi random graph model and prove that the subsequent classification can be asymptotically optimal, outperforming the entire graph under high-dimensional conditions. Our method is evaluated on various simulated datasets and real-world human and murine graphs derived from functional and structural magnetic resonance images. The results demonstrate its excellent performance in estimating the ground-truth signal subgraph and achieving superior classification accuracy.
keywords:
iterative screening, distance correlation, graph classification
@@ 1 Introduction
The analysis of graph structure is critical in various big data fields, including neuroscience, internet mapping, and social networks [23, 21, 3, 39, 34, 14]. Due to the large size of graphs in practice, such as in social networks and raw neuroimages, it is often necessary to use smaller subgraphs from the observed graphs. Moreover, the selected subgraph should maintain or improve subsequent inference. For example, identifying a subset of brain regions from brain imaging to better predict the phenotype of interest in each subject.
The statistical problem of feature reduction and dimension selection has been extensively studied, with well-known methods such as Lasso [37], adaptive Lasso [44], Dantzig selector [5], sure independence screening [10, 17], among others. These methods have specific objectives, such as sparsity and the recovery of ground-truth. Among them, the screening method is known for its computational efficiency and model-free nature and are commonly used for high-dimensional data [43], making it a suitable candidate for graph data.
However, dimension reduction for graph data presents unique challenges because of its high-dimensionality and the unique structure of the $n\times n$ adjacency matrix. To that end, this paper proposes an iterative screening method on graph data, which utilizes distance-based correlation and independence screening in an iterative manner to estimate the signal subgraph. During each iteration, we define the feature for each vertex using the adjacency of the reduced graphs, compute a distance-based correlation between the feature and the label of interest $Y$, and discard vertices with small correlations. This process is repeated recursively on the reduced graphs from previous iterations, yielding a smaller set of vertices each time until an estimated signal subgraph is selected for output.
@@ 2.1 Setting and Notations
In the classical statistical pattern recognition setting, the pairs of observations $\{(A_{i},Y_{i})\}_{i=1}^{m}$ are independent and identically distributed pairs according to a distribution $F_{A,Y}$ [8], that is
@@ 2.2 Distance Correlation
The distance correlation is a measure that can detect all types of dependencies between two random variables, given sufficient sample size [36]. To compute the sample distance correlation, two pairwise distance matrices are transformed and multiplied using a Hadamard product. The sample distance correlation is asymptotically 00 if and only if the two underlying random variables are independent. For more detailed mathematical information about the distance correlation and its population definition, see the Appendix.
The distance correlation is a computationally efficient method [30], and has been shown to be equivalent to kernel correlation [33]. It has been used for various inference tasks [40, 11, 29, 28, 24] not limited to screening. This paper also utilizes a local version of the distance correlation called the multiscale graph correlation (MGC), which improves testing power against nonlinear dependencies [38, 32, 16, 27].
@@ 4 Theoretical Properties
  • The graph adjacency matrix $A$ follows the Erdos-Renyi random graph model [9] and has a bounded Frobenius norm.
@@
The results suggest that using the estimated signal subgraph via iterative screening can be expected to perform better than using the whole graph, when the size of the signal subgraph is fixed and the number of observed graphs is comparable to the size of the whole graph. This setting is illustrated in the top panel of Figure 2 with $n=200$, $m=300$, and $|S|=20$, and is also observed in the experiments in Section 5 where the estimated subgraph yields better classification performance. All proofs and additional mathematical background are provided in the appendix.
@@ 5.1 Signal Subgraph Estimation
We estimate the subgraph using various screening methods, including the conventional screening with Dcor and MGC, iterative screening with Dcor and MGC at $\delta=0.5$ and $\delta=0.05$ respectively, and screening with canonical correlation analysis (CCA) [13] and RV coefficient (RV) [26]. Since the actual size of the signal subgraph was known to be 20, we output the estimated subgraph at the same size and calculated the true positive rate. The ROC curve is shown in Figure 1, while Table 1 reports the AUC and runtime for each approach. We observe that Dcor and MGC outperform CCA and RV, and terative screening improves the performance over conventional screening. Furthermore, iterative screening with $\delta=0.05$ yields better results than iterative screening with $\delta=0.5$ at the cost of a longer running time.
@@ 6.1 Site and Sex Prediction With Human Brain
Our objective is to predict the sex and site of each individual based on functional magnetic resonance image (fMRI) graphs [22]. We utilized two datasets, SWU4 [18] and HNU1 [6], which include $467$ and $300$ subjects, respectively. Each individual’s fMRI scan is registered to the MNI152 template using the Desikan atlas, which has $70$ regions [7]. The graphs are created using the NeuroData’s MRI Graphs pipeline [^1] https://github.com/neurodata/ndmg, a popular tool for processing and representing brain images.
@@ 6.2 Sex Difference in Mouse Brain
Structural magnetic resonance imaging has provided insight into the genetic basis of mouse brain variability by examining the relationship between volume covariance and genotypes [2]. With high-resolution diffusion tensor imaging and tractography, we can now investigate the underlying bases for structural connectivity patterns [4], in relationship with genotype and sex. Based on MRI and conventional Nissl histology, we scanned and registered $55$ mouse brains (pooled genotypes) into the space of a minimum deformation template, aligned to Waxholm space [15]. The atlas labels were propagated onto the template and, subsequently, onto each individual brain using ANTs [1]. We employed DSI Studio [41] to estimate tract-based structural connectivity for each brain, which was then represented as a graph with $332$ vertices, $166$ per hemisphere. Of the $55$ mice, $32$ are male, and $23$ are female.
Similarly, we conduct a leave-one-out evaluation using an iterative vertex screening to estimate the signal subgraph, followed by a $9$-nearest-neighbor classifier to predict the left-out sample based on the estimated signal subgraph. Figure 6 demonstrates the prediction error and distance correlation when using the iterative screening algorithm. Despite the small sample size and fluctuating prediction error, the screening method outputs a signal subgraph of size $10$, which results in a near-optimal classification error of $0.18$. The estimated signal vertices include a thalamic component and the periaqueductal gray, which play an important role in driving sexually dimorphic mouse brain development [35, 25].
@@ 7 Conclusion
In summary, we developed an iterative vertex screening methodology to estimate the signal subgraph of interest and successfully applied the method in simulations and real data. Utilizing distance correlation and multiscale graph correlation lends strong interpretability to our methods. Given the existence of signal vertices where each vertex is dependent on the graph-level attributes (while non-signal vertices are independent), Theorem 1 suggests that the signal subgraph can be recovered with probability converging to $1$ as the number of graphs increases. Furthermore, Theorems 2 and 3 suggest that subsequent classification using the signal vertices can be asymptotically Bayes optimal, and in certain cases (depending on the relationship between $n$ and $m$), better than utilizing the full graph.
We shall emphasize that the proposed method is essentially a dimension reduction technique, and one could use any subsequent classifier, not just Bayes plug-in. Therefore, Theorems 2 and 3 should be viewed as providing theoretical guarantees under a simple classifier case. Nevertheless, Theorem 1 is a very general result that focuses solely on dimension reduction. Intuitively, excluding independent vertices is expected to benefit many subsequent tasks beyond classification. Previous research has demonstrated that dependence measures can lead to better interpretability and improvements in complex machine learning architectures [12, 42].
The experiments and theories provide strong evidence that the iterative approach effectively and accurately estimates the signal subgraph, resulting in better performance for subsequent classification compared to conventional screening methods. It is important to emphasize once again that our method requires multiple graph datasets with a common set of vertices, and it has been shown that incorporating more graphs can enhance subsequent vertex classification [31]. If there are multiple graphs available but the vertices are not matched, then our method is not applicable. However, if part of the vertex set is matched across graphs, our method can still be applied to the matched vertex subset. In cases where the vertices are matched but the actual correspondence is unknown, graph matching techniques may be applied first [20, 19].
@@ Acknowledgement
A.1 Technical Preliminaries
A.1.1 Distance Correlation
Given sufficient sample size, the distance correlation [36] is able to detect all types of dependencies between two random variables. The population distance covariance $Dcov(X,Y)$ can be defined via either the characteristic functions or Euclidean distance as:
@@ A.1.1 Distance Correlation
A.1.2 Graph Classification
We introduce the binary classification setting of predicting the label $Y\in\{0,1\}$ using graph $A$. This set-up serves as the basis for Section 4 and the simulations. The network model under consideration is the inhomogeneous Erdos-Renyi (ER) random graph model [9], which allows edges to have different probabilities and generates a family of distributions on undirected graphs. The ER model can also be viewed as a stochastic block model with each block containing only one vertex.
@@ A.1.2 Graph Classification
Given this model, the optimal classification performance is achieved by the Bayes classifier $g^{*}(\cdot)$ [8] defined as
@@
A.2 Theorem Proof
A.2.1 Proof of Theorem 1
@@
We have now met the two requirements to apply Theorem 1 in [17]. By utilizing the theorem and choosing $\kappa=0$, we can conclude that there exist two positive constants $c_{1},c_{2}$, and for any $20<\gamma<1/$ 2 we have
@@ A.2.2 Proof of Theorem 2
Setting $2(e_{V}\epsilon_{2}+\epsilon_{1})=\epsilon$ and $2\epsilon_{1}^{2}=c_{4}\epsilon_{2}^{2}$, we have $\epsilon_{2}=\frac{\epsilon}{2e_{V}+\sqrt{2c_{4}}}$. Applying Theorem 2.3 in [8] yields